How to calculate ND1 and ND2
1
Expert's answer
2020-07-28T14:51:59-0400

ND1 and ND2 are associated with the Black-Scholes model for option pricing. The model is used to determine the fair or theoretical value of a put or a call option.


Firstly, D1 and D2 are first calculated according to the formulae, and a rounded to 3 decimal places.


Secondly, ND1 and ND2 are calculated from the standardized cumulative normal distribution table (the Z - tables).


Suppose, D1 = 0.077, then to find ND1 one will reference a standardized normal distribution table. Under z column we look for 0.0, then proceed to look for 7 under the middle column and this will give 0.5279. The final digit (7) is looked under the final column, there is 28 under 7. Now, to get ND1 = N(0.077) = 0.5279 + 0.0028 = 0.5307


The same applies for ND2. Suppose D2 = 0.098

ND2 = N(0.098)

Looking for 0.0 under z column and for 9 under the middle column will give 0.5359. By looking under 8 in the last column, the value 32 is obtained. Therefore, ND2 = N(0.098) = 0.5359 + 0.0032

= 0.5391


The table is provided below,